### v1

# When in Doubt, Deliberate: Confidence-Based Routing to Expert Debate for Sexism Detection

Abstract
Sexist content online increasingly appears in subtle, context-dependent forms that evade traditional detection methods. Its interpretation often depends on overlapping linguistic, psychological, legal, and cultural dimensions, which produce mixed and sometimes contradictory signals, even in annotated datasets. These inconsistencies, combined with label scarcity and class imbalance, result in unstable decision boundaries and cause fine-tuned models to overlook subtler, underrepresented forms of harm. Together, these limitations point to the need for a design that explicitly addresses the combined effects of (i) underrepresentation, (ii) noise, and (iii) conceptual ambiguity in both data and model predictions. To address these challenges, we propose a two-stage framework that unifies (i) targeted training procedures to adapt supervision to scarce and noisy data with (ii) selective, reasoning-based inference to handle ambiguous or borderline cases. Our training setup applies class-balanced focal loss, class-aware batching, and post-hoc threshold calibration to mitigate label imbalance and noisy supervision. At inference time, a dynamic routing mechanism classifies high-confidence cases directly and escalates uncertain instances to a novel Collaborative Expert Judgment (CEJ) module, which prompts multiple personas and consolidates their reasoning through a judge model.Our approach achieves state-of-the-art results across several benchmarks, with a +2.72% improvement in F1 on the EXIST 2025 Task 1.1, and a gains of +4.48% and +1.30% on the EDOS Tasks A and B, respectively.
When in Doubt, Deliberate: Confidence-Based Routing to Expert Debate for Sexism Detection
Anwar Alajmi1,2, Gabriele Pergola1 1Department of Computer Science, University of Warwick, Coventry CV4 7AL, UK 2College of Business Studies, Public Authority of Applied Education and Training, Kuwait {anwar.alajmi, gabriele.pergola.1}@warwick.ac.uk
11footnotetext: Warning: This paper includes examples that may contain explicit, offensive, or harmful language. 22footnotetext: Code will be released upon acceptance.
1 Introduction
Figure 1: Overview of the proposed routing mechanism. Predictions with low confidence from the fine-tuned model are escalated to a Collaborative Expert Judgment (CEJ) module, where multiple LLM-based personas engage in structured debate, before the final decision.
Sexist content on social media is increasingly pervasive, often appearing in subtle, sarcastic, or context-dependent forms that evade traditional classification methods. From seemingly innocuous stereotypes to overt hate speech, such content contributes to the emotional and psychological harm that disproportionally affects women abercrombie-etal-2023-resources.
These diverse forms of sexism are shaped by distinct yet overlapping factors: linguistic, psychological, legal, and cultural, among others. As a result, both human annotators and automated systems must contend with mixed and sometimes contradictory signals: a message may appear harmless to a general audience, yet be perceived as harmful by a legal expert, a psychologist, or someone with lived experience of gender-based discrimination davani2022disagreement.
This multidimensionality creates several challenges for annotation and automatic detection. Even the largest and most reliable datasets edos23; exist25 are typically small, highly imbalanced, and marked by substantial annotator disagreement not only due to ambiguity in individual examples, but also because different annotators may implicitly prioritize different criteria (e.g., linguistic form, psychological impact, cultural normativity, or legal definitions) khan2025explaining. These mixed signals can overshadow the representation of less explicit but equally damaging forms of sexism, introducing noise and instability into both the data and the models. At the same time, large language models (LLMs) are pretrained on vast, heterogeneous data sources, absorbing social norms, stereotypes, which are inconsistently represented, and at times in contradiction with human feedback. As a result, fine-tuning these large models on small, noisy sexism datasets is often insufficient to resolve these conflicts, and may inadvertently reinforce dominant or explicit patterns while neglecting subtler cases khan2025explaining. These limitations underscore the need for models that explicitly account for the combined effect of underrepresentation, noise, and conceptual ambiguity in both data and predictions.
To address these challenges, we introduce a two-stage framework that combines (i) targeted training procedures to adapt supervision to scarce and noisy data, with (ii) selective, reasoning-based inference to handle ambiguous cases. This joint design enables the specialised model to make efficient decisions when possible, while engaging with finer contextual reasoning when necessary. During (i) training, we mitigate long-tailed label distributions and inconsistent annotation through three well-known training strategies. These have never been adapted into an integrated framework for such tasks: the class-balanced focal loss (CB-Focal) (ce), which amplifies the contribution of minority and difficult examples, and the class-aware batching, which ensures uniform exposure to underrepresented classes during optimization. We further employ post-hoc threshold calibration to stabilize the decision boundaries of the specialist classifier and improve robustness under noisy supervision. At (ii) inference time, we deploy confidence-aware dynamic routing, which filters the model predictions by difficulty: high-confidence samples are directly classified, while uncertain cases are escalated to a Collaborative Expert Judgment (CEJ) module. The CEJ module implements a structured reasoning pipeline composed of multiple personas, such as a linguist, a gender expert, a psychologist, and an everyday speaker, each prompted to analyze the same input from distinct perspectives. Their deliberations are subsequently synthesized by a judge model, which integrates these perspectives into a final, interpretable decision (Figure 1). The aim is not to faithfully replicate human interaction, but rather to approximate its structure to activate the reasoning and knowledge that instruction-tuned LLMs already encode, yet rarely employ in single-step classification, as evidenced by the limited predictive power of zero-shot baselines edos23; exist25; khan2025explaining.
Our experimental evaluation on both the EDOS and EXIST 2025 datasets demonstrates that our methods outperform existing state-of-the-art approaches across almost all tasks. On EXIST 2025 (Task 1.1), the combination of targeted training procedures and confidence-aware routing yields an increase of over 5 F1 points, while also setting new benchmarks on ICM and ICM-Norm metrics. For EDOS, we observe significant improvements in both binary (Task A) and multi-class (Task B) classification: selective escalation with the CEJ module boosts macro F1 by 3.9 points on Task A and  1 point on Task B. Notably, our approach surpasses previous baselines without employing any ensemble or data augmentation methods.
Our contributions can be summarized as follow:
  • A unified framework with selective reasoning capability: We introduce a unified framework for sexism detection that combines the unmatched efficiency of neural classifiers and targeted training procedures, such as class-balanced focal loss, class-aware batching, and post-hoc threshold calibration, with a reasoning-based inference mechanism that is selectively applied to challenging cases.
  • Collaborative Expert Judgment: We introduce a novel module, Collaborative Expert Judgment (CEJ), that leverages multi-persona prompting to resolve ambiguous or borderline cases, enabling the activation of reasoning encoded in instruction-tuned LLMs.
  • Comprehensive experimental assessment: We provide a comprehensive analysis of where multi-persona reasoning delivers the greatest benefit, particularly in handling ambiguous and minority cases, and discuss the limitations under severe class imbalance. We demonstrate that integrating these components achieves state-of-the-art results on almost all tasks of both the EXIST 2025 and EDOS datasets.
2 Related Work
Our work is aligned to at least two research line on sexissm detection and persona-based analyses.
Sexism Classification.
Annotator disagreement, often shaped by cultural perspectives, complicates sexism detection davani2022disagreement. Initiatives like the EXIST 2024 shared task encouraged methods robust to inconsistent annotations. Successful approaches leveraged transformer-based models such as DistilBERT sanh2019distilbert, DeBERTa he2020deberta, RoBERTa liu2019roberta, and hybrid ensembles that utilized data augmentation and multi-task learning. When it comes to LLMs, previous studies have employed adversarial evaluations and data augmentation to enhance detection accuracy. For example, samory2021call highlighted standard model vulnerabilities through psychologically-grounded adversarial datasets, emphasizing the need for psychologically informed robustness. Similarly, khan2025explaining addressed annotator disagreements through definition-driven data augmentation and ensemble methods, improving data-level robustness.
Beyond explicit abuse, a growing challenge lies in detecting nuanced or implicit forms of hate. The work by zeng2025sheeps show that LLMs often fail to recognize metaphorical expressions of hate speech, highlighting substantial gaps in current moderation capabilities. Their findings emphasize that even state-of-the-art models struggle when hateful intent is obscured through indirect language, a phenomenon that is particularly relevant for sexist discourse. Building on this, the authors in fasching2025model demonstrate that hate speech detection outcomes vary significantly across LLM-based systems, underscoring how moderation decisions remain model-dependent and inconsistent. Our work directly responds to these challenges by integrating confidence-aware routing and multi-expert arbitration, ensuring that inconsistent or ambiguous cases are addressed through structured reasoning and synthesis.
Persona-based Approaches.
More recent work integrates multi-agent architectures, expert simulations, and human-in-the-loop reasoning to improve transparency and social norm alignment. While expert persona-driven prompting strategies, such as those in xu2023expertprompting; long2024multi, have advanced interpretability, our method surpasses static personas by introducing dynamic, structured, and confidence-aware interactions among multiple expert roles, capturing richer contextual nuances. Moreover, our framework integrates a dedicated judge model to arbitrate conflicts, providing nuanced decisions beyond simple voting mechanisms.
Emerging research demonstrated the effectiveness of multi-expert deliberation in LLM-based decision-making. lu-etal-2024-triageagent introduce a multi-agent framework for clinical triage where LLM agents assume distinct roles, engage in multi-round discussions with self-confidence scoring, and iteratively revise their assessments until reaching consensus. Similarly, long2024multi simulate multiple domain experts within a single LLM, aggregate their responses, and select the best one. Both approaches show that structured discussion among diverse perspectives yields more reliable outputs than single-viewpoint reasoning, as cross-perspective validation mitigates individual biases and enables principled conflict resolution before final judgment. Our framework extends this paradigm to sexism detection, where the inherently subjective nature of the task makes perspectival diversity essential.
Another notable study that effectively leverages expert agents is Expert-Token-Routing chai2024expert, which adopts a router-based approach similar to that of jiang2023llm. In this method, a router model is trained to direct each question or instruction to a specialized expert, where the expert LLM is responsible for generating the next token. Their approach encodes expert LLMs as special tokens within the vocabulary of the meta-LLM, drawing inspiration from ToolkenGPT hao2023toolkengpt.
Our progressive prompt refinement aligns with iterative human-in-the-loop methods discussed in shah2025prompt, systematically enhancing logical coherence and prompting effectiveness. In addition, we preserve the logical structure and enhance the prompt formatting for effective model reasoning as it is shown by li2025llms’s experiments.
Figure 2: CEJ framework, where a collaborative reasoning discussion among personas with refinement is generated before the judge model makes a definitive decision.
3 Methodology
We introduce a two-stage framework for sexism detection. The framework combines a (i) specialized neural model with (ii) a selective reasoning component. The specialized model is fine-tuned using adapted methodologies to improve robustness under both data imbalance and noisy supervision. At inference, a confidence-aware routing mechanism selects high-confidence instances for the specialised model or escalates uncertain ones to a Collaborative Expert Judgment (CEJ) module, where multiple personas deliberate and a judge model produces the final decision (Figure 2).
3.1 Domain-Tuned Model
We fine-tune a small instruction-tuned LLM (LLaMA-3.2-3B) as a domain-specialized classifier. To handle severe class imbalance, we integrate three complementary strategies: Class-Balanced Cross-Entropy (CB-CE) (ce) to reweight loss by effective sample count, Focal Loss (ce) to emphasize difficult minority-class instances, and Class-Aware Batch Sampling (CAB) (henning23) to ensure balanced gradient contributions. For binary classification, we apply temperature scaling (guo2017calibration) and tune a decision threshold on the development set; for multi-class classification, we use CB-Focal loss given the more severe imbalance.
3.2 Confidence-Aware Routing
At this stage, all the low-confidence predictions from the specialised models are escalated to the collaborative reasoning method for a multi-perspective analysis. Specifically, we adopt a selective classification rule (elyaniv2010foundations) given an input $x$, the model produces a predicted label $\hat{y}_{s}$ with the associated confidence score $c_{s}(x)$:
$\hat{y}_{s}$ $=\arg\max_{k}p_{\theta}(y{=}k\mid x)$ (1)
$c_{s}(x)$ $=\max_{k}p_{\theta}(y{=}k\mid x)$ (2)
If $c_{s}(x)\geq\tau_{\text{conf}}$, we accept $\hat{y}_{s}$; otherwise, the input is routed to the CEJ module.
In multi-class settings, low confidence may not fully capture uncertainty, particularly when the model is uncertain between its top two predictions. We define the margin between the two most probable labels as $m(x)=p_{(1)}(x)-p_{(2)}(x)$ where $p_{(1)}(x)$ and $p_{(2)}(x)$ denote the largest and second-largest posterior probabilities. An instance is then routed to CEJ if:
$\big(c_{s}(x)<\tau_{\text{conf}}\big)\;\land\;\big(m(x)<\tau_{\text{margin}}\big)$ (3)
3.3 Collaborative Expert Judgment (CEJ)
Prior research on abusive and sexist language highlights the importance of combining expert knowledge with lived experience sap2020socialbiasframes; davani2022disagreement. Inspired by this, we define a set of personas representing diverse interpretive perspectives. Expert roles, such as linguists, psychologists, gender studies scholars, and legal professionals have been shown to provide critical perspectives on language use, structural bias, emotional impact, and legal norms lazar2005feminist; vidgen2019challenges. To still capture non-expert and experiential perspectives, we also incorporate a sexism victim persona, reflecting direct experience of harm, and a layperson persona, capturing everyday interpretations that frequently underlie annotator disagreement (Table A3).
It is important to emphasize that the aim of this multi-persona prompting is not to simulate faithful and exhaustive expert panels, but rather to elicit and organize the contextual, normative, and experiential knowledge already encoded within large language models. By inducing the model to articulate these diverse perspectives, the CEJ module encourages the application of domain understanding that often remains latent in single-step classification.
# Variant Model ICM ICM-Norm F1_YES F1 Zero-shot Baselines 1 LLaMA-3.3-70B-Instruct 0.3002 0.6509 0.6868 0.7145 2 Cogito-70B 0.3656 0.6837 0.7109 0.7348 3 Qwen2.5-72B-Instruct 0.3765 0.6892 0.7112 0.7384 4 Cogito-70B (reasoning) 0.3772 0.6896 0.7143 0.7384 EXIST 2025 Task 1.1 Leaderboard 5 BERT nowakowski2025exist 0.5727 0.7878 0.7802 6 XLM-RoBERTa pan2025exist 0.5799 0.7915 0.7824 7 Dual-Transformer Fusion Network khan2025exist 0.5806 0.7918 0.7837 8 DeepSeek-R1-Distill-Llama-8B villarreal2025exist 0.6127 0.8079 0.7945 9 Ensemble Approach c:22 0.6249 0.8141 0.7991 10 XLM-RoBERTa c:21 0.6297 0.8165 0.7996 11 LLaMA-3.1-8B-Instruct tian2025marioexist2025simple 0.6774 0.8405 0.8167 Ours 12 $\mathcal{C}_{1}$ LLaMA-3.2-3B 0.5709 0.7913 0.7596 0.7934 13 $\mathcal{C}_{2}$ $\mathcal{C}_{1}$ + Task-specific optimizations 0.6595 0.8315 0.8089 0.8257 14 $\mathcal{C}_{3}$ $\mathcal{C}_{2}$ + Routing to CEJQ→C 0.6831 0.8433 0.8191 0.8348 15 $\mathcal{C}_{4}$ $\mathcal{C}_{2}$ + Routing to CEJL→C 0.6963 0.8500 0.8233 0.8389
Table 1: Performance on the EXIST 2025 Task 1.1 test set. First, zero-shot baselines with generalist LLMs; then, leaderboard submissions at EXIST 2025 Task 1.1; finally, our routing pipeline. “Q” and “L” denote the models used for personas debate (Qwen2.5-72B-Instruct and LLaMA-3.3-70B, respectively), and “C” refers to the judge model Cogito-70B (reasoning mode). Bold indicates the best score, and underline denotes the runner-up.
3.3.1 Structured Debate and Judge
We operationalize the CEJ module as a multi-stage panel, in which each persona offers a distinct interpretive perspective on the input. The process comprises four sequential stages, designed to elicit, interrogate, and synthesize domain-relevant knowledge:
Initial Opinions. Each persona independently analyzes the input, providing an initial classification, justification, with a confidence score. This initial stage establishes diverse baseline perspectives grounded in each persona’s unique expertise.
Structured Debate. Personas are subsequently exposed to all initial opinions and engage in critical evaluation of their peers’ reasoning. During this deliberation, each persona must: (i) engage with at least one other perspective, explicitly agreeing or disagreeing and offering supporting rationale; (ii) revise their own stance if confronted with compelling counter-arguments; and (iii) reassess the input through the interpretive lens of other personas, re-evaluating both intent and target. This process results in a revised classification, an updated justification, and an adjusted confidence score for each persona.
Summarization. A dedicated summarization agent condense the full debate into a concise synthesis highlighting the main arguments, points of consensus, and unresolved disagreements.
Final Judgment. Finally, the judge model synthesizes all available evidence, i.e., the original input, both initial and revised persona opinions, and the debate summary, to produce a final adjudication. Prompted to act impartially, the judge delivers a definitive label, comprehensive justification, and confidence score that resolves any lingering disagreements.
# Variant Model Task A Task B Task C Zero-shot Baselines 1 Qwen-2.5-72B-Instruct 0.6190 0.4246 0.2576 2 LLaMA-3.3-70B-Instruct 0.6421 0.4297 0.2775 3 Cogito-70B 0.6106 0.4214 0.2683 4 Cogito-70B (reasoning) 0.6332 0.4325 0.2826 SemEval 2023 Task 10.A Leaderboard 5 DeBERTa-v3-large + twHIN-BERT-large zhou-2023-pinganlifeinsurance 0.8746 6 RoBERTa-Large + ELECTRA 0.8740 0.7203 0.5487 7 DeBERTa Ensemble 0.8740 Data Augmentation and Ensemble Methods 8 SEFM zhong-etal-2023-uirisc 0.8538 0.6619 0.4641 9 QCon feely-etal-2023-qcon 0.8400 0.6400 0.4700 10 HULAT segura-bedmar-2023-hulat-semeval 0.8298 0.5877 0.4458 11 CSE khan2025explaining 0.8819 0.7243 0.5639 12 DDA khan2025explaining 0.8769 0.7277 0.6018 13 PaLM Ensemble 0.7326 Ours 14 $\mathcal{C}_{1}$ LLaMA-3.2-3B 0.7824 0.5941 0.3717 15 $\mathcal{C}_{2}$ $\mathcal{C}_{1}$ + Task-specific optimizations 0.8986 0.6279 0.4189 16 $\mathcal{C}_{3}$ $\mathcal{C}_{2}$ + Routing to CEJQ→C 0.9195 0.7324 0.5842 17 $\mathcal{C}_{4}$ $\mathcal{C}_{2}$ + Routing to CEJL→C 0.9214 0.7421 0.5904
Table 2: Macro-F1 scores on the EDOS datasets with zero-shot baselines, leaderboard references, augmentation/ensemble methods, and our approach. Bold = best score, underline = runner-up.
4 Experiments
This section describes the datasets, evaluation metrics, experimental setup, and results used to assess the proposed framework.
4.1 Experimental Setup
Datasets. We evaluate our methods on two widely used benchmark corpora for sexism detection: the EXIST 2025 Tweets dataset exist25 and the EDOS dataset edos23 (detailed descriptions in Appendix A.1). The EXIST dataset is selected for its difficulty, as it incorporates annually updated tweets, capturing emerging sexist language, evolving slang, and annotation disagreements that mirror genuine ambiguities in human judgment. The EDOS dataset offers a complementary perspective with high quality fine-grained annotations at multiple levels of granularity, organized across three hierarchical classification tasks.
Evaluation Metrics. Following the official setup, we use the Evaluate ALL 2.0 tool evall2025 for scoring under the EXIST 2025 Task 1.1 (Hard-Hard) configuration. The evaluation metrics include: (i) Inter-Consistency Measure (ICM) amigo2022evaluating, which quantifies agreement between predicted and reference annotations by combining intra-system consistency and penalizing overlap; (ii) ICM-Norm, a normalized version of ICM scaled to $[0,1]$ for comparability; and the (iii) F1 score, the harmonic mean of precision and recall. As for the EDOS dataset, we follow the official evaluation protocol and report results using the Macro-F1 score, the unweighted mean of the F1-scores computed independently for each class across Task A, B, and C (Appendix A.1).
Baselines. We evaluate our proposed framework against both fine-tuned and zero-shot LLM baselines. These focuses on state-of-the-art approaches for sexism detection; while an exhaustive exploration of all available LLMs is beyond the scope of this work. As specialized models, we employe LLaMA-3.2-3B, adapted via LoRA fine-tuning (hu2021lora) on both the EDOS and EXIST training datasets. Full task-specific training information are available in Appendix A.2. To evaluate instruction-following and reasoning abilities in a zero-shot setting, we test Qwen2.5-72B-Instruct qwen2025qwen25technicalreport, LLaMA-3.3-70B-Instruct grattafiori2024llama, and Cogito-70B deepcogito_cogito_v2 using simple classification prompts (Appendix A.3). Pre-trained checkpoints were obtained from HuggingFace huggingface. All LLM experiments were executed locally through LangChain langchain and Ollama ollama on a system equipped with three Nvidia A100 GPUs.
4.2 Results
In this section, we present the experimental results of our framework and baselines on the EXIST 2025 and EDOS benchmarks.
EXIST Dataset. Results for the the EXIST 2025 Task 1.1 test set, binary classification, are reported in Table 1. The first group of models comprises LLMs evaluated in a zero-shot setting. Among these, Cogito-70B (reasoning mode) achieves the highest F1_YES (0.7143), with closely matched performance from Qwen2.5-72B-Instruct (0.7112) and standard Cogito-70B (0.7109). Overall, all zero-shot models with basic prompts underperform compared to leaderboard approaches, reflecting the limitations of out-of-the-box LLMs for specialized task. The second group reports official leaderboard systems submitted to EXIST 2025 Task 1.1. Here, LLaMA-3.1-8B-Instruct tian2025marioexist2025simple leads in both ICM (0.6774) and F1_YES (0.8167). Transformer fusion architectures and BERT-based systems round out the group, showing the persistence of competitive results of earlier architectures.
The final section of Table 1 details the results of our routing framework. We build upon the LLaMA-3.2-3B speciliased model which achieves performance on par with lower-ranking leaderboard models (F1_YES: 0.7596). Applying targeted optimizations in $\mathcal{C}_{2}$ (multi-dataset training, class-balanced loss, threshold calibration) yields substantial improvements: 0.8089 F1_YES and 0.8257 macro-F1, with results that already surpass many leaderboard submissions. Further gains are realized through confidence-aware dynamic routing. Configuration $\mathcal{C}_{3}$ integrates Qwen2.5-72B-Instruct for persona-based debate, reaching 0.8191 F1_YES, while $\mathcal{C}_{4}$, leveraging LLaMA-3.3-70B for personas and Cogito-70B as judge, achieves the highest scores across all reported metrics. This configuration outperforms the best leaderboard system by 0.66 points on F1_YES and establishes a new benchmark for the task.
EDOS Dataset. We evaluate the results on the EDOS dataset, reporting the macro-F1 across three classification tasks of increasing granularity (Table 2). Results are presented for zero-shot LLMs, official leaderboard systems, models employing data augmentation and ensembling, and our selective routing framework.
The first group report the results of instruction-tuned LLMs evaluated in a zero-shot setting. Performance remains limited across all tasks when using basic prompts. The best macro-F1 for binary classification (Task A) is achieved by LLaMA-3.3-70B-Instruct (0.6421), closely followed by Cogito-70B (reasoning mode, 0.6332) and Qwen2.5-72B-Instruct (0.6190). For fine-grained classification (Task C), all zero-shot models fall below 0.29. Among official leaderboard systems from SemEval 2023 Task 10.A, ensemble methods based on DeBERTa and RoBERTa architectures set a strong baseline, reflecting the effectiveness of large pre-trained transformer models with supervised fine-tuning.
The next group encompasses systems employing data augmentation, model ensembling, or both. Here, CSE khan2025explaining and DDA khan2025explaining achieve notable gains: CSE reaches 0.8819 on Task A and 0.5693 on Task C, while DDA records the highest reported macro-F1 on Task C (0.6018) and competitive results on other tasks.
Our approach in the base configuration, $\mathcal{C}_{1}$ (fine-tuned LLaMA-3.2-3B), achieves competitive results for binary and multi-class tasks, but falls short on fine-grained Task C (0.3717). Introducing targeted training optimizations in $\mathcal{C}_{2}$, including multi-dataset training, class-balanced loss, and threshold calibration, yields marked improvements, reaching 0.8986 on Task A, 0.6279 on Task B, and 0.4189 on Task C. Finally, confidence-aware escalation to the CEJ module provides substantial further gains. Configuration $\mathcal{C}_{3}$, which employs Qwen2.5-72B-Instruct for persona-based debate, achieves 0.9195 on Task A and 0.7324 on Task B. The strongest results are obtained by $\mathcal{C}_{4}$, where LLaMA-3.3-70B is used for the persona debate and Cogito-70B as the judge, achieving 0.9214 on Task A and 0.7421 on Task B. Both configurations outperform all previous systems on binary and multi-class classification. On Task C, both routing variants remain competitive (0.5904 and 0.5842), though they do not surpass the best-performing data augmentation method (DDA). This gap can be attributed to the cascading effect of baseline limitations. Despite employing task-specific optimization techniques, ($\mathcal{C}_{2}$) exhibits progressive macro-F1 score degradation as class granularity increases (0.8986 $\rightarrow$ 0.6279 $\rightarrow$ 0.4189 across Tasks A, B, and C), indicating that these optimization techniques only partially mitigate the challenge of learning 11 fine-grained categories with severe class imbalances. Since CEJ routes instances based on $\mathcal{C}_{2}$’s confidence scores, this weaker starting point propagates through the pipeline. In contrast, DDA addresses data scarcity directly by synthesizing training examples that populate underrepresented categories, providing a stronger foundation that inference-level refinement alone cannot replicate.
Category n $\mathcal{C}_{2}$ $\mathcal{C}_{3}$ $\mathcal{C}_{4}$ Gain
Task B
1. Threats/Harm 89 .681 .848 .857 +17.6
2. Derogation 454 .727 .608 .745 +1.9
3. Animosity 333 .563 .663 .519 +10.1
4. Prejudiced Disc. 94 .541 .811 .848 +30.7
Task C
1.1 16 .217 .300 .250 +8.3
1.2 73 .626 .771 .767 +14.5
2.1 205 .568 .670 .646 +10.2
2.2 192 .537 .571 .561 +3.4
2.3 57 .426 .514 .528 +10.2
3.1 182 .700 .739 .728 +3.9
3.2 119 .556 .735 .724 +18.0
3.3 18 .231 .973 .973 +74.2
3.4 14 .100 .933 .933 +83.3
4.1 21 .408 .069 .188 -22.1
4.2 73 .238 .152 .198 -4.1
Table 3: Class-wise F1 on EDOS Tasks B and C. $\mathcal{C}_{2}$: domain-tuned model; $\mathcal{C}_{3}$: CEJ routing (Qwen); $\mathcal{C}_{4}$: CEJ routing (LLaMA). Gain: improvement from $\mathcal{C}_{2}$ to best CEJ configuration (percentage points).
Table 3 presents class-wise F1 scores for Tasks B and C. CEJ shows an inverse relationship between class frequency and effectiveness: rarer categories benefit most, with gains of +30.7 (Prejudiced Discussions, nn=94) and +17.6 (Threats/Harm, nn=89) in Task B, while the most frequent category (Derogation, nn=454) improves only marginally (+1.9). This pattern holds in Task C, where the smallest classes 3.3 (nn=18) and 3.4 (nn=14) achieve gains of +74.2 and +83.3 points. This suggests that multi-perspective debate compensates for representation bias in underrepresented categories.
However, CEJ degrades performance for categories 4.1 (-22.1) and 4.2 (-4.1). These subcategories, 4.1 (Supporting mistreatment of individual women) and 4.2 (Supporting systemic discrimination), require subtle distinctions: 4.1 separates victim-blaming from neutral risk statements, while 4.2 separates endorsement of discrimination from good-faith debate about the progress of women’s rights. The domain-tuned model learns these annotation-specific boundaries directly from training data. CEJ, however, introduces multiple interpretive perspectives that may not align with these boundaries, adding variance where the task requires consistency with annotator conventions.
Figure 3 presents per-persona gains after the debate phase. The legal studies persona shows the largest improvement (+13.1% ICM, +4.1% F1), as legal assessments inherently require synthesizing intent, context, and harm—dimensions enriched through cross-persona deliberation. Psychologist (+8.7%) and gender studies expert (+7.9%) also benefit substantially.
Figure 3: Performance gain after the debate phase.
Ablation Study
Table 4 presents the prompt ablation results. The full prompt configuration ($\mathcal{P}_{5}$) consistently outperforms ablated variants across all personas. The largest degradations occur when removing sexism definitions ($\mathcal{P}_{2}$) or persona-specific context ($\mathcal{P}_{1}$), with the Psychologist exhibiting the greatest sensitivity (+19.3% ICM gain from $\mathcal{P}_{1}$ to $\mathcal{P}_{5}$). Notably, ICM and ICM-Norm improvements exceed F1 gains, indicating enhanced alignment with human annotation distributions rather than mere accuracy.
Persona $\mathcal{P}_{1}$ $\mathcal{P}_{2}$ $\mathcal{P}_{3}$ $\mathcal{P}_{4}$ $\mathcal{P}_{5}$
Normal Person .442 .459 .561 .593 .598
Linguist .439 .446 .584 .563 .586
Gender Studies Expert .433 .443 .577 .576 .583
Legal Studies Expert .466 .453 .558 .566 .573
Sexism Victim .433 .469 .568 .571 .592
Psychologist .406 .446 .574 .593 .599
Table 4: Prompt ablation (ICM scores, LLaMA-3.3-70B). $\mathcal{P}_{1}$: baseline few-shot prompt with role identity. $\mathcal{P}_{2}$: expands the role into a specialised expert persona. $\mathcal{P}_{3}$: introduces formal sexism definitions aligned with annotation guidelines. $\mathcal{P}_{4}$: adds nuanced and multilingual examples. $\mathcal{P}_{5}$: incorporates refined guidelines informed by error analysis.
5 Conclusion
In this work we addressed some of the structural issues in sexism detection, including data scarcity, class imbalance, and annotation noise; factors that often undermine the reliability of standard fine-tuning. We proposed a unified framework that augments targeted training with confidence-calibrated routing and multi-persona deliberation. Through a combination of robust loss functions, balanced batching, and post-hoc calibration, our system improves specialist reliability, while selectively escalating ambiguous cases to a structured, multi-perspective judgment module, CEJ. Evaluations on EXIST 2025 and EDOS show that this approach not only advances state-of-the-art performance, but also proves especially effective on cases marked by ambiguity or underrepresentation.
Limitations
While our frameworks demonstrate strong performance and interpretability through expert-guided multi-persona reasoning, several limitations remain. Our approaches utilize prompt engineering and LLM-generated content, which can reflect biases present in the models’ pretraining data. Additionally, the scalability of multi-agent interactions poses a practical constraint: as the number of simulated personas increases, so does computational cost, which may hinder applicability in real-time or resource-limited settings.
Moreover, our system’s performance degrades with highly granular classification schemes. The domain-tuned baseline deteriorates with increasing class granularity and severe class imbalances, and this limitation propagates through the routing pipeline. This suggests that reasoning-based refinement is better suited for binary or moderate-granularity tasks, while extreme class fragmentation with heavy imbalances may require complementary data-level interventions such as the DDA method. Importantly, these approaches are not mutually exclusive; CEJ could potentially be combined with data augmentation strategies to leverage both enriched training distributions and structured inference-time reasoning. We leave exploration of such hybrid architectures to future work.
Appendix A Appendix
A.1 Datasets Details and Task Structure
The EDOS dataset (SemEval-2023 Task 10) consists of 14,000 training, 2,000 development, and 4,000 test examples for Subtask A (binary sexism detection) including 970 sexist examples for Subtasks B and C. It defines a three-level hierarchical classification problem:
  • Task A: Binary classification — distinguishing between sexist and not sexist texts.
  • Task B: Four-way categorization of sexist content, assigning each sexist instance to one of the following categories:
    1. 1.
      Threats, plans to harm and incitement
    2. 2.
      Derogation
    3. 3.
      Animosity
    4. 4.
      Prejudiced discussions
  • Task C: Fine-grained subcategorization, where each sexist instance is further assigned to one of the following subcategories:
    1. 1.1
      Threats of harm
    2. 1.2
      Incitement and encouragement of harm
    3. 2.1
      Descriptive attacks
    4. 2.2
      Aggressive and emotive attacks
    5. 2.3
      Dehumanising attacks & overt sexual objectification
    6. 3.1
      Casual use of gendered slurs, profanities, and insults
    7. 3.2
      Immutable gender differences and gender stereotypes
    8. 3.3
      Backhanded gendered compliments
    9. 3.4
      Condescending explanations or unwelcome advice
    10. 4.1
      Supporting mistreatment of individual women
    11. 4.2
      Supporting systemic discrimination against women as a group
For Tasks B and C, we train on filtered subsets where label_sexist = “Sexist” from Task A predictions. Each task uses a separate specialist model with task-specific label mappings.
In addition, we evaluate on the EXIST 2025 dataset, which focuses on sexism detection in a cross-lingual, cross-platform setting. It contains 10,034 binary-labeled samples (YES or NO) in English and Spanish. The dataset includes 6,920 training samples (3,260 English, 3,660 Spanish), 1,038 development samples (489 English, 549 Spanish), and 2,076 test samples (978 English, 1,098 Spanish). The task is structured as follows:
  • Task 1.1 (EXIST): Binary classification of social media posts as Sexist vs. Not Sexist. This task is analogous to EDOS Task A, but differs in scale, language diversity (English and Spanish), and domain coverage (Twitter and Gab), thereby presenting a more challenging generalization setting.
A.2 Training Details
A.2.1 Shared Training Infrastructure Details (Binary and Multi-Class)
We employ LLaMA-3.2-3B as our base model with 4-bit quantization to enable efficient fine-tuning on consumer hardware while maintaining model quality. The quantization configuration is as follows:
  • Quantization type: NF4 (Normal Float 4-bit)
  • Double quantization: Enabled
  • Compute dtype: bfloat16 (if supported) or float16
LoRA adapters are applied to all attention and feed-forward projection layers (q_proj, k_proj, v_proj, o_proj, gate_proj, up_proj, down_proj). Configuration varies by task (see Table A1). This reduces trainable parameters to $1\%$ of the full model while maintaining competitive performance across all tasks.
A.2.2 Binary Classification Optimizations
Class-balanced cross-entropy (CB-CE) is emplyed to address class imbalance by re-weighting loss contributions based on effective sample counts. For each class $y$ with $n_{y}$ training examples, the effective number accounts for information overlap:
$\text{EN}_{y}=\frac{1-\beta^{n_{y}}}{1-\beta},\quad\text{where }\beta=0.999$ (4)
Raw class weights are computed as $w_{y}^{\text{raw}}=1/\text{EN}_{y}$. These are normalized to unit mean:
$w_{y}^{\text{norm}}=w_{y}^{\text{raw}}\cdot\frac{C}{\sum_{c=1}^{C}w_{c}^{\text{raw}}}$ (5)
To prevent destabilization from extreme weight ratios, we apply clamping:
$w_{y}=\text{clip}(w_{y}^{\text{norm}},w_{\min},w_{\max})$ (6)
with $w_{\min}=0.25$ and $w_{\max}=4.0$, followed by re-normalization to unit mean. Hence the loss function is formulated as follows:
$\mathcal{L}_{\text{CB-CE}}(x,y)=-w_{y}\sum_{c=1}^{C}\tilde{y}_{c}\log\text{softmax}(\mathbf{z}(x))_{c}$ (7)
Temperature Scaling
After training, we calibrate model confidence using temperature scaling to improve probability estimates. We fit a single temperature parameter $T$ by minimizing negative log-likelihood on the development set:
$\begin{split}T^{\star}&=\arg\min_{T}\sum_{(x,y)\in\mathcal{D}_{\mathrm{dev}}}\Bigg[-\frac{z_{y}(x)}{T}\\ &\qquad\qquad\quad+\log\!\sum_{c=1}^{C}\exp\!\left(\frac{z_{c}(x)}{T}\right)\Bigg]\end{split}$ (8)
Calibrated probabilities are then computed as:
$p_{\theta}^{T}(y\mid x)=\text{softmax}(\mathbf{z}(x)/T^{\star})_{y}$ (9)
Temperature scaling is particularly important for binary classification as it corrects systematic over-confidence or under-confidence in model predictions without requiring retraining.
Decision Threshold Tuning
Threshold $t$ is tuned on calibrated probabilities to maximize development set macro-F1:
$t^{\star}=\operatorname*{arg\,max}_{t\in[0,1]}\;\mathrm{F1}_{\mathrm{macro}}\big(\hat{\mathbf{y}}_{t}\big),\;\hat{\mathbf{y}}_{t}\!=\!\big(\hat{y}_{t}(x)\big)_{x\in\mathcal{D}_{\mathrm{dev}}}$ (10)
This approach balances precision and recall across both classes, particularly important when minority class performance is critical. In practice, $t^{\star}$ typically falls in the range $[0.3,0.6]$ depending on class imbalance severity.
A.2.3 Multi-Class Classification Optimizations
For multi-class settings with long-tailed distributions (Tasks B and C), we adopt Class-Balanced Focal Loss, which combines effective-number weighting with focal loss modulation. The CB-Focal loss for a sample $(x,y)$ is:
$\mathcal{L}_{\text{CB-Focal}}(x,y)=-w_{y}(1-p_{y})^{\gamma}\log(p_{y})$ (11)
where:
  • $w_{y}$ is the CB weight for class $y$ (computed as in the previous section)
  • $p_{y}$ is the predicted probability for the true class
  • $\gamma$ is the focusing parameter (typically $\gamma\in[1,3]$, we use $\gamma=2$)
The $(1-p_{y})^{\gamma}$ term down-weights easy examples (high $p_{y}$) while amplifying hard or misclassified ones (low $p_{y}$). This dual mechanism:
  • Re-weights by class frequency (via $w_{y}$)
  • Re-weights by example difficulty (via $(1-p_{y})^{\gamma}$)
This loss is particularly effective for Tasks B and C, where rare subcategories (e.g., backhanded compliments, condescending explanations) are overwhelmed by frequent classes (derogation). By combining class re-weighting and hard-example mining, CB-Focal improves macro-F1 by 3–5 points, ensuring minority classes are not neglected while reducing over-confidence on majority classes.
Class-Aware Batching.
Standard uniform sampling produces batches dominated by majority classes (e.g., 12–14 majority, 2–4 minority instances), leading to inefficient learning on minority classes. Given per-device batch size $B$ and $C$ classes, the quota per class is:
$k=\lfloor B/C\rfloor$ (12)
For Task B with $C=4$ classes and $B=16$, we sample $k=4$ instances per class per batch. For each training batch:
  1. 1.
    Partition training data into class-specific subsets $\mathcal{D}_{1},\ldots,\mathcal{D}_{C}$
  2. 2.
    For each class $c\in\{1,\ldots,C\}$, randomly sample $k$ instances from $\mathcal{D}_{c}$ with replacement
  3. 3.
    Concatenate samples to form batch $\mathcal{B}$ of size $k\cdot C\leq B$
  4. 4.
    Shuffle combined batch to avoid positional biases
This ensures minority classes contribute equally despite smaller pool sizes. Without replacement would exhaust small classes quickly, leading to epoch-boundary effects.
A.2.4 Training Hyperparameters
Table A1 summarizes the hyperparameters selected across tasks.
Hyperparameter Task 1.1/Task A Task B Task C
Learning rate $2\times 10^{-4}$ $6\times 10^{-5}$ $2\times 10^{-5}$
Batch size 16 16 16
Gradient accumulation 2 2 2
Training epochs 5 8 12
Warmup ratio 0.1 0.1 0.1
Weight decay 0.01 0.01 0.01
Max sequence length 512 512 512
Label smoothing ($\epsilon$) 0.05 0.05 0.05
CB-CE $\beta$ 0.999 0.999 0.999
CB-CE $w_{\min}$ 0.25 0.25 0.25
CB-CE $w_{\max}$ 4.0 4.0 4.0
Focal loss $\gamma$ 2.0 2.0
LoRA rank ($r$) 32 96 96
LoRA alpha ($\alpha$) 64 192 192
LoRA dropout 0.1 0.2 0.2
Table A1: Hyperparameter configuration across tasks.
A.3 Zero-shot Baselines
For the zero-shot baselines, LLMs including Qwen2.5-72B-Instruct, LLaMA-3.3-70B-Instruct, and Cogito-70B were evaluated using task-specific simple prompts. Each model was queried with a zero-shot classification instruction tailored to the corresponding EDOS task as follows:
  • EXIST Task 1.1/ EDOS Task A: “Classify the following text as sexist or not sexist.”
  • Task B: “Classify the following text into one of the sexism categories: 1) Threats, plans to harm and incitement; 2) Derogation; 3) Animosity; 4) Prejudiced discussions.”
  • Task C: “Classify the following text into one of the sexism subcategories: 1.1) Threats of harm; 1.2) Incitement and encouragement of harm; 2.1) Descriptive attacks; 2.2) Aggressive and emotive attacks; 2.3) Dehumanising attacks & overt sexual objectification; 3.1) Casual use of gendered slurs, profanities, and insults; 3.2) Immutable gender differences and gender stereotypes; 3.3) Backhanded gendered compliments; 3.4) Condescending explanations or unwelcome advice; 4.1) Supporting mistreatment of individual women; 4.2) Supporting systemic discrimination against women as a group.”
A.4 Personas Progressive Prompt Design
Our progressive prompt construction method incrementally integrates role-conditioning, reasoning mechanisms, expert knowledge, and guideline refinements into our prompt design approach. Each iteration aims at improving performance, interpretability, and alignment with human annotator behavior as listed below:
  • $\mathcal{P}_{1}$: A basic prompt with few-shot examples and role identity. This sets the task framing and label expectations, testing the model’s generalization capabilities.
  • $\mathcal{P}_{2}$: Enhances reasoning by expanding the role identity and background into a specialized expert persona. This stage encourages more structured reasoning by clearly separating the task into sub-steps,
  • $\mathcal{P}_{3}$: Integrates formal sexism definition aligned with annotation guidelines.
  • $\mathcal{P}_{4}$: Adds nuanced and multilingual examples to strengthen generalization.
  • $\mathcal{P}_{5}$: Based on manual error analysis, the prompt incorporates refined guidelines, including considerations of the tweet author’s intent and target audience, as well as the inclusion of edge cases (e.g., profanity and slang) to guide borderline decisions (examples provided in Table A2).
# Example
$\mathcal{P}_{1}$ You are a psychologist. Classify the tweets as sexist or not sexist. Here are some examples…
$\mathcal{P}_{2}$ + (45 y/o female, Argentina, Ph.D. Psychology) Focus on language, cognitive biases, and the psychological effects of sexism. (1) Analyze the text carefully; (2) Think before responding; (3) Classify the tweet as sexist (1) or not sexist (0).
$\mathcal{P}_{3}$ + Sexism is gender-based prejudice, stereotyping, or discrimination, typically against women. Label as sexist (1) if the tweet: (a) is sexist itself, (b) describes a sexist situation, or (c) criticizes sexist behavior.
$\mathcal{P}_{4}$ + "She got promoted because they needed ’more women in leadership.’" → Sexist (1).
"No tengo nada contra las mujeres, pero en cargos altos siempre rinden menos." → Sexist (1).
$\mathcal{P}_{5}$ + (1) Consider the author’s intent (insult, joke, venting, shaming). (2) Profanity alone does not indicate sexism — consider the context carefully.
Table A2: Examples illustrating the structure and progression of each prompt stage.
The selected personas are prompted with unique descriptions as shown in Table A3. The six expert personas used in our classification framework are described, each representing a distinct interpretive lens on sexism. These roles were carefully constructed to reflect diverse perspectives, including legal, psychological, linguistic, and lived experience. To increase realism, demographic and educational attributes were heuristically aligned with annotator metadata from the EXIST 2025 dataset, mimicking the diversity and subjectivity found in real-world human annotation.
Persona Perspective Role Description (Prompt) Layperson General public perspective An average person (18–22 y/o male, Portugal, B.A.) with no specialized expertise in law, psychology, or linguistics. Linguist Language patterns and bias A linguist (23–45 y/o male, Poland, B.A. Linguistics) specializing in semantics, pragmatics, and discourse analysis, with a focus on gendered language. Psychologist Emotional and cognitive effects A psychologist specializing in language, cognitive biases, and the psychological effects of sexism. Legal Expert Legal view on gender-based discrimination A legal expert (46+ y/o male, Portugal, M.A. Law) specializing in anti-discrimination laws, workplace regulations, and gender equality. Gender Expert Structural and ideological analysis A gender studies expert (46+ y/o female, UK, B.A. Gender Studies) with deep knowledge of gender theories, power dynamics, and social structures. Sexism Victim Personal impact of experiencing sexism A person (18–22 y/o female, South Africa, H.S. diploma) who has personally experienced sexism and understands its emotional and social impact.
Table A3: Persona perspectives and detailed role descriptions.
A.4.1 Preliminary Investigation
Table A4 reports a preliminary investigation conducted on relatively smaller-scale LLMs (7–8B parameters), namely Mistral-7B, Dolphin3-8B, andCogito-7B. The table compares the baseline prompt (“Classify the given text as sexist or not sexist”) against our first enhanced persona-based prompt, denoted as $\mathcal{P}_{1}$, evaluated on the EXIST 2025 development set. In contrast to the baseline, $\mathcal{P}_{1}$ incorporates few-shot examples and an explicit role identity (e.g., psychologist, linguist, normal person) to guide model reasoning.
Overall, the results confirm that even modest prompt enhancements yield consistent improvements across metrics. For Mistral-7B andCogito-7B, $\mathcal{P}_{1}$ produced reliable gains in ICM, normalized ICM, and F1, while Dolphin3-8B, despite its weak baseline performance, exhibited the largest relative improvements. Importantly, the absolute best scores are consistently achieved byCogito-7B, suggesting that stronger base models are less sensitive to prompt design but still benefit incrementally. In contrast, weaker models (e.g., Dolphin3-8B) depend more heavily on the additional structure introduced by persona framing.
Finally, performance varied across personas: roles such as Normal Person and Legal Studies Expert saw more pronounced gains, while Sexism Victim and Linguist produced only modest improvements. These exploratory findings were not used as final results but rather as guidance for subsequent large-scale experiments, helping us refine prompt design choices and identify promising personas for sexism detection.
Persona Model ICM1 ICM2 ICM%↑ Norm1 Norm2 Norm%↑ F11 F12 F1%↑
Psychologist Mistral-7B 0.2388 0.4059 +69.96% 0.6195 0.7031 +13.49% 0.7433 0.8019 +7.89%
Dolphin3-8B -0.2299 -0.2299 0.00% 0.3850 0.3850 0.00% 0.5245 0.5245 0.00%
Cogito-7B 0.4230 0.4703 +11.16% 0.7116 0.7353 +3.33% 0.8052 0.8233 +2.25%
Sexism Victim Mistral-7B -0.0258 0.3559 +1478.68% 0.4871 0.6781 +39.16% 0.6672 0.7882 +18.13%
Dolphin3-8B -0.2265 -0.1298 +42.70% 0.3867 0.4351 +12.51% 0.5244 0.5866 +11.86%
Cogito-7B 0.4089 0.4505 +10.16% 0.7046 0.7254 +2.95% 0.8024 0.8148 +1.54%
Linguist Mistral-7B 0.1882 0.3435 +82.56% 0.5941 0.6718 +13.11% 0.7327 0.7800 +6.44%
Dolphin3-8B -0.3184 -0.3184 0.00% 0.3407 0.3407 0.00% 0.4724 0.4724 0.00%
Cogito-7B 0.3967 0.4055 +2.22% 0.6985 0.7029 +0.63% 0.7957 0.8011 +0.68%
Legal Studies Expert Mistral-7B 0.2317 0.3586 +54.75% 0.6159 0.6794 +10.30% 0.7494 0.7858 +4.86%
Dolphin3-8B -0.2300 0.0427 +118.57% 0.3849 0.5214 +35.45% 0.5261 0.6693 +27.26%
Cogito-7B 0.4375 0.4772 +9.08% 0.7189 0.7387 +2.75% 0.8117 0.8254 +1.69%
Normal Person Mistral-7B 0.3445 0.3952 +14.72% 0.6723 0.6977 +3.78% 0.7816 0.7985 +2.16%
Dolphin3-8B -0.2395 -0.0675 +71.82% 0.3802 0.4662 +22.62% 0.5212 0.6213 +19.20%
Cogito-7B 0.3982 0.4863 +22.11% 0.6992 0.7433 +6.31% 0.7978 0.8287 +3.87%
Gender Studies Expert Mistral-7B 0.2875 0.3956 +37.57% 0.6438 0.6979 +8.41% 0.7590 0.7986 +5.22%
Dolphin3-8B -0.2493 -0.0769 +69.15% 0.3753 0.4615 +22.99% 0.5188 0.6164 +18.78%
Cogito-7B 0.3826 0.4339 +13.41% 0.6914 0.7171 +3.72% 0.7930 0.8098 +2.12%
Table A4: Preliminary results of persona-based prompts, comparing baseline prompt with $\mathcal{P}_{1}$. Metrics include ICM, normalized ICM, and F1, with relative improvements (%↑).
A.5 Collaborative Expert Judgment (CEJ) Algorithm and Complexity
Algorithm 1 summarizes the CEJ pipeline for a dataset of $N$ instances and a persona set of size $K$. CEJ requires $K$ persona inferences, one discussion, one summarization, and one judgment per instance, yielding $N\times(K+3)$ total LLM calls with complexity $O(NK)$. The cost therefore scales linearly with both the dataset size and the number of personas.  [^2] The analysis below concerns the interaction-level complexity of the CEJ procedure rather than low-level runtime or token usage.
Algorithm 1 Collaborative Expert Judgment (CEJ)
Input: Dataset $\mathcal{D}=\{(x_{i},\text{id}_{i})\}_{i=1}^{N}$, persona set $\mathcal{Y}=\{y_{1},\dots,y_{K}\}$, Large Language Model $\mathcal{M}$
Output: Opinion matrix $\mathcal{O}=\{O_{i,k}\}$, summary $S_{i}$, and final judgment $J_{i}$ for each $x_{i}$
1:for each instance $(x_{i},\text{id}_{i})\in\mathcal{D}$ do
2:  Initial Persona Classification
3:  for each persona $y_{k}\in\mathcal{Y}$ do
4:   Construct persona-specific classification prompt $\mathcal{P}_{\text{cls}}(x_{i},y_{k})$
5:   $O_{i,k}\leftarrow\mathcal{M}.\texttt{invoke}(\mathcal{P}_{\text{cls}})$
6:  end for
7:  Simulated Panel Discussion
8:  Construct discussion prompt $\mathcal{P}_{\text{disc}}(x_{i},\{O_{i,k}\}_{k=1}^{K},\mathcal{Y})$
9:  $\mathcal{D}_{i}\leftarrow\mathcal{M}.\texttt{invoke}(\mathcal{P}_{\text{disc}})$
10:  Summary Generation
11:  Construct summarization prompt $\mathcal{P}_{\text{sum}}(\mathcal{D}_{i},\mathcal{Y})$
12:  $S_{i}\leftarrow\mathcal{M}.\texttt{invoke}(\mathcal{P}_{\text{sum}})$
13:  Final Judgment with Justification
14:  Construct judgment prompt $\mathcal{P}_{\text{judge}}(x_{i},\{O_{i,k}\}_{k=1}^{K},S_{i})$
15:  $J_{i}\leftarrow\mathcal{M}.\texttt{invoke}(\mathcal{P}_{\text{judge}})$
16:end for
A.6 Qualitative Analysis
Figures A1 and A2 illustrate the Collaborative Expert Judgment (CEJ) framework using a representative tweet from EXIST 2025. Figure A1 shows the initial classification stage, where six expert personas independently assess the input. Each provides a binary label (YES/NO), a domain-grounded justification, and a confidence score. Most personas classify the tweet as sexist, citing trivialization of female emotional expression or reinforcement of gendered stereotypes; only the Legal Expert disagrees, arguing the statement lacks discriminatory intent under legal criteria.
Figure A2 presents the structured debate stage, where personas reflect on others’ reasoning and may revise their stance and confidence. This deliberative process demonstrates how LLM-based agents engage in structured argumentation—aligning with or challenging one another based on shared evidence and domain-specific reasoning. While most personas maintain their original YES classification, their justifications become more nuanced and confidence scores are recalibrated. The Legal Expert again maintains NO, citing insufficient legal grounds.
Figure A3 illustrates successful error correction via confidence-based routing. Here, the specialist misclassifies the input as YES (sexist) with confidence below threshold $\tau$, triggering CEJ escalation. The expert panel classifies it as NO, reasoning that the text targets physical appearance rather than gender, lacks patriarchal power dynamics, and occurs in a casual context. The judge concurs, overriding the false positive. This case demonstrates how routing captures borderline errors: low-confidence predictions are delegated to multi-perspective deliberation, where richer contextual and intent analysis corrects specialist mistakes.
Figure A1: An example output from the initial classification stage of the CEJ framework.
Figure A2: An example output from the structured debate stage of the CEJ framework.
Figure A3: Error correction via confidence-based routing. The specialist’s false positive (confidence $\tau$) is overridden by CEJ deliberation.
Appendix B Prompts
This section presents the prompt templates used throughout In the CEJ framework. Initial persona judgments are collected using the prompt shown in Figure A4. This is followed by the structured debate prompt (Figure A5), which enables each persona to reflect on peer opinions, revise its stance if persuaded, and recalibrate its confidence. The discussion is then summarized and passed to the judge prompt (Figure A6), which concludes the process by generating a final label along with a justification and calibrated confidence score. All prompts rely on consistent definitions and structured output formats (e.g., JSON objects with fields such as label, justification, and confidence) to support interpretability and downstream evaluation.
You are {personas_description} tasked with classifying the following tweet for sexism. Tweet: {tweet_text} Your task: 1. Read the guidelines below carefully 2. Analyze the tweet carefully for sexism. 3. Think before responding. 4. Decide the final label: 1 (sexist) or 0 (not sexist). 5. Provide a short justification for their label based on their role. 6. Output a confidence score between 0.0 and 1.0 reflecting your certainty. Sexism Definition {definition}
Objective:
{objective}
Here are some examples:
{examples}
Output Example: Provide only a valid JSON object like the following example:
{ "persona": "Normal Person", "label": "1", "justification": "The tweet stereotypes womens intelligence.", "confidence": "0.87" }
Figure A4: Template for initial persona classification in the CEJ framework.
You are continuing the expert panel discussion on the following tweet: Tweet: {tweet_text} Initial Opinions: {persona_opinions} Now, each persona must: 1. Read all other personas’ initial opinions. 2. Reflect on whether their own reasoning is still the strongest. 3. Engage with at least one other persona by agreeing or disagreeing with their argument. 4. Update their stance if persuaded, or affirm their original decision. 5. Reassess and adjust their confidence accordingly. Sexism Definition {definition}
Objective:
{objective}
Here are some examples:
{examples}
Important Notes:
Confidence can be increased if supported by solid reasoning, or reduced if uncertainty arises. Final answers must state if the stance is changed or unchanged. Output Example (per persona): { "persona": "Sexism Victim", "intent": "The authors intent is to shame the woman by dismissing her distress.", "reaction": "Agree with Linguist because their interpretation highlights the use of gendered stereotypes.", "updated_reasoning": "While my initial view focused on tone, I now realize the tweet uses the victim card trope to discredit womens emotional responses.", "final_stance": 1 (changed from 0), "updated_confidence": 0.72 }
Figure A5: Structured discussion prompt for reasoning and stance revision in the CEJ framework.
You are an impartial judge of the detection of sexist language. Review the following expert panel summary and the original tweet being discussed then classify it for sexism. You have the following information: Tweet: {tweet_text} Discussion summary:
{summary}
Your task: 1. Read the guidelines below carefully 2. Review expert panel summary. 3. Analyze the tweet for sexism. 4. Think before responding. 5. Decide the final label: 1 (sexist) or 0 (not sexist). 6. Provide a brief justification for your decision. 7. State your confidence score (0.0–1.0). Sexism Definition: {definition}
Objective:
{objective}
Here are some examples:
{examples}
Output Example: Provide only a valid JSON object like the following:
{ "label": 0, "justification": "Although the tweet contains profanity and strong language, it is not clearly directed at women. Based on the personas disagreement and the classification guidelines, the tweet is offensive but not sexist.", "confidence": 0.79 }
Figure A6: Prompt for the judge revision following persona discussion in the CEJ method.